OpenStudy (heyitslizzy13):

can anyone explain the process to me? I have an idea but I just can't seem to get the answer

OpenStudy (heyitslizzy13):

h

OpenStudy (taramgrant0543664):

1 league is equal to 3 nautical miles This is a 1:3 ratio So if you have 610 leagues you would have x nautical miles To find your x you would just multiply 610 by 3 to get it into nautical miles

OpenStudy (anonymous):

Another way I like to think about finding conversion factors is just find out what's equal, so like @taramgrant0543664 said \[1 \text{ league} = 3 \text{ nautical miles}\] So to create our conversion factor we first recognize that we can multiply anything by 1 and it doesn't change it. So we divide this equation appropriately: \[1 = \frac{3 \text{ nautical miles}}{1\text{ league} }\] Why? Well because of this! \[610 \text{ leagues} * 1 = 610 \text{ leagues} * \frac{3 \text{ nautical miles}}{1\text{ league} }\] Here I just plugged in this fancy form of 1. Remember how multiplying by one doesn't change anything? It doesn't change the actual distance, just our representation of it. Now here's the trick: \[\frac{\text{leagues}}{\text{leagues}}=1\] Right? Anything divided by itself is just 1! So we use that to simplify this here: \[610 \text{ leagues} * \frac{3 \text{ nautical miles}}{1\text{ league} } = 610 * 3 \text{ nautical miles}\] And there you go, you're a multiplication away from your final answer. :D